OpenStudy (anonymous):

A sphere having a 10 cm radius has a maximum possible error in measurement of 0.02 cm. If you want to find the volume of this sphere, what will be the maximum error in the calculated volume?

OpenStudy (anonymous):

The maximum error is 0.02cm so the maximum radius is 10.02cm. From this: \[4/3*\pi(10.02)^{3}-4/3*\pi(10)^{3}\] \[4/3*\pi(10.02^{3}-10^{3})=25.18cm^{3}\]